Your firm buys a new equipment for Rs. 60,000. It will have a useful life of six years, with zero salvage value. It is to be depreciated by the SL method over six years, with placement at the start of the first year. It will give your firm savings of Rs. 16,000 per year. The tax rate will be 13%. However, you decide to sell it at the end of the third year, for Rs. 28,000. Calculate the after-tax IRR for this?


1
Expert's answer
2021-01-08T12:57:47-0500



"NPV=\sum\frac{CF}{1+r}^t-IC"


"IRR=\frac{NPV}{IC}\times 100"

it turns out very low IRR it is not profitable to sell equipment